# MMRole: A Comprehensive Framework for Developing and Evaluating Multimodal Role-Playing Agents

### v1

@@ 6.5 Evaluation Results and Analyses
As shown in Table 5, we report the average results across all test samples for each evaluated MRPA, along with the detailed results on both the in-distribution test set (In-Test) and the out-of-distribution test set (Out-Test) for our MMRole-Agent. Notably, although some of the scores of MRPAs exceed $1$, this does not necessarily mean that their performance is superior to that of GPT-4, which we use to construct MMRole-Data. To generate multi-turn dialogue data, we use a single GPT-4 API call to produce the entire dialogue directly. In contrast, when testing MRPAs, we supply dialogue histories and require MRPAs to generate responses. This method is relatively easier, but it is challenging to ensure the consistency of multi-turn dialogues if using this approach for data construction.
In the MRPA group with over 100 billion parameters, Claude 3 Opus exhibits superior performance. Meanwhile, in the MRPA group with tens of billions of parameters, LLaVA-NeXT-34B achieves the highest performance. Finally, in the MRPA group with billions of parameters, MMRole-Agent is the best. Notably, LLaVA-NeXT-34B outperforms Gemini Pro Vision, while LLaVA-NeXT-Mistral-7B surpasses Yi-VL-34B. This demonstrates that both the training methods and training data are equally important for enhancing LMMs, rather than merely expanding the model size.
@@
As depicted in Figure 3, we provide a clear visual representation of the evaluation results. The overall performance ranking of MRPAs closely aligns with their specific performance ranking on each metric. However, significant differences exist in the score variations across various metrics. Specifically, all MRPAs achieve high scores on the Fluency metric with minimal variations, suggesting that producing fluent content is not a challenge for current LMMs. Conversely, there are notable differences among MRPAs on other metrics, particularly Personality Consistency and Tone Consistency. This reveals that multimodal understanding abilities and role-playing qualities are more challenging aspects that require attention in the development of MRPAs.
@@ 7 Conclusion
In this paper, we propose the concept of Multimodal Role-Playing Agents (MRPAs) for the first time by extending RPAs with multimodal understanding abilities. Moreover, we construct MMRole-Data, a large-scale, high-quality dataset for developing and evaluating MRPAs. To stably and comprehensively assess MRPAs, we introduce MMRole-Eval, a robust evaluation method using a specialized reward model, comprising eight metrics across three dimensions. Comprehensive evaluation results indicate that MMRole-Agent, the first specialized MRPA, exhibits improved performance and strong generalization capabilities. Additionally, multimodal understanding abilities and role-playing qualities are more challenging aspects that require attention in the development of MRPAs.
Appendix A List of Characters
Figure 4 lists all character constructed in MMRole-Data, with the series to which the characters in the out-of-distribution test set belong being underlined.
Appendix B Examples of Character Profiles
@@
Appendix C Examples of the Two-Stage Generation Process for Hypothetical Real-Life Characters
Figure 7 and Figure 8 presents the prompts used to generate meta information and expand detailed profiles for hypothetical real-life characters, whereas Figure 9 exemplifies meta information for five randomly selected hypothetical real-life characters.
Appendix D Prompts for Dialogue Generation
@@
Appendix E Prompts for Querying MRPAs
Appendix F Prompts for GPT-4 Scoring
Appendix G Prompts for Reward Model Scoring

### v2

@@ 6.5 Evaluation Results and Analyses
As shown in Table 5, we report the average results across all test samples for each evaluated MRPA, along with the detailed results on both the in-distribution test set (In-Test) and the out-of-distribution test set (Out-Test) for our MMRole-Agent. Notably, although some of the scores of MRPAs exceed $1$, this does not necessarily mean that their performance is superior to that of GPT-4, which we use to construct MMRole-Data. To generate multi-turn dialogue data, we use a single GPT-4 API call to produce the entire dialogue directly. In contrast, when testing MRPAs, we supply dialogue histories and require MRPAs to generate responses. This approach is relatively easier, but it is challenging to ensure the consistency of multi-turn dialogues if used for data construction.
In the MRPA group with over 100 billion parameters, Claude 3 Opus exhibits superior performance. Meanwhile, in the MRPA group with tens of billions of parameters, LLaVA-NeXT-34B achieves the highest performance. Finally, in the MRPA group with billions of parameters, MMRole-Agent is the best. Notably, LLaVA-NeXT-34B outperforms Gemini Pro Vision, while LLaVA-NeXT-Mistral-7B and MMRole-Agent surpass Yi-VL-34B. This suggests that both the training methods and training data are important for enhancing LMMs, rather than merely expanding the model size.
@@
As shown in Figure 3, we provide a clear visual representation of the evaluation results. The overall performance rankings of MRPAs closely align with their specific rankings on each metric. However, significant differences exist in the score variations across various metrics. Specifically, all MRPAs achieve high scores on the Fluency metric with minimal variations, suggesting that producing fluent content is not a major challenge for current LMMs. Conversely, there are notable differences among MRPAs on other metrics, particularly on Personality Consistency and Tone Consistency. It reveals that multimodal understanding abilities and role-playing qualities are more challenging aspects that require attention in the development of MRPAs.
@@
Table 6: The average results across all test samples for each evaluated RPAs. ‘w/o vision’ signifies that image information is excluded from the input prompt of RPAs.
RPAs Comment. Human-Role. Inter-Role. Overall
GPT-4 Turbo w/o vision 0.5746 1.1330 1.0843 0.9306
GPT-4 Turbo 1.0261 1.2275 1.3450 1.1995
Claude 3 Opus w/o vision 0.3290 1.1803 1.1420 0.8838
Claude 3 Opus 1.0088 1.2889 1.3916 1.2298
MMRole-Agent w/o vision 0.4192 0.8909 0.7907 0.7003
MMRole-Agent 1.0450 0.9556 0.9619 0.9875
Additionally, to highlight the inherent advantages of MRPAs over single-modal RPAs, we conduct comparative experiments on two SOTA general-purpose LMMs and our MMRole-Agent. As presented in Table 6, we report the Image-Text Relevance scores on the Out-Test set evaluated by GPT-4, where ‘w/o vision’ signifies that image information is excluded from the input prompt of RPAs. The results clearly demonstrate that excluding image information significantly reduces the Image-Text Relevance of all RPAs’ responses, particularly in commentary interaction scenarios. In multi-turn human-role and inter-role dialogue scenarios, textual dialogue history could provide indirect clues about images, leading to relatively smaller declines in the Image-Text Relevance scores compared to those in commentary interactions. Nevertheless, the absence of visual inputs still results in a marked drop in performance across all scenarios.
Table 7: The results on the Out-Test set for MMRole-Agent with different numbers of characters.
Characters Number of Characters Overall
The Avengers 16 0.965
The Avengers + Other English Fictional Characters 34 0.968
The Avengers + Hypothetical Real-Life Characters 36 0.970
ALL (ours) 72 0.983
Table 8: The average results across all test samples for MMRole-Agent with different numbers of training samples.
Training Data Number of Samples Overall
SAMPLE 8.5K 0.967
ALL (ours) 85K 0.989
6.6 Detailed Analyses of MMRole-Agent
To demonstrate the superiority of our MMRole-Agent, use cases of MMRole-Agent, GPT-4, and QWen-VL-Chat are presented and analyzed in Appendix K. Moreover, sensitivity test results detailed in Appendix L indicate that MMRole-Agent is compatible with different prompts and does not exhibit signs of overfitting. The strong performance and generalization abilities of our MMRole-Agent can be primarily ascribed to the following two factors:
  1. 1.
    Training with Large-Scale, High-Quality Data: The training set of MMRole-Data comprises 72 characters, 11K images, and over 85K samples. Furthermore, as depicted in Figure 1(a) and Figure2, due to the well-designed data construction pipeline, meticulous manual annotation and quality control, and the use of GPT-4, the data is of high quality. This large-scale, high-quality dataset enables MMRole-Agent to comprehensively learn the instruction demands, knowledge, and abilities in multimodal role-playing. To verify this point, we compare the performance differences between a model trained on the full dataset (ALL) and a model trained on a randomly sampled subset consisting of one-tenth of the data (SAMPLE), both evaluated after one epoch of training. As shown in Table 8, the performance of ALL is superior to that of SAMPLE.
  2. 2.
    Joint Training with Diverse Multi-Character Data: We incorporate data from 72 diverse characters to jointly train a unified MMRole-Agent. This approach, akin to the principles of multi-task learning, enables the model to acquire generalizable multimodal role-playing capabilities, rather than being confined to specific characters. To verify this point, we first train a model using data of characters from The Avengers, then gradually add additional characters to the training set for subsequent models. As presented in Table 7, we evaluate the performance of each model on the Out-Test set. The models’ zero-shot performance steadily improves as more characters are incorporated. Notably, with comparable numbers of characters, introducing hypothetical real-life characters (with significant differences from The Avengers) yields greater gains than adding other English fictional characters, indicating the significance of training with diverse data.
@@ 7 Conclusion
In this paper, we propose the concept of Multimodal Role-Playing Agents (MRPAs) for the first time by extending RPAs with multimodal understanding abilities. Moreover, we construct MMRole-Data, a large-scale, high-quality dataset for developing and evaluating MRPAs. To stably and comprehensively assess MRPAs, we introduce MMRole-Eval, a robust evaluation approach that comprises eight metrics across three dimensions, scoring MRPAs with the ground truth for comparison by a specialized reward model. Evaluation results reveal that our MMRole-Agent, the first specialized MRPA, exhibits improved performance and strong generalization capabilities. Additionally, multimodal understanding abilities and role-playing qualities are more challenging aspects that require attention in the development of MRPAs. However, there exists a limitation that the training data for MMRole-Agent is primarily synthesized by GPT-4, which constrains its performance from surpassing GPT-4 itself. In future work, we will address this limitation by leveraging multiple SOTA LMMs respectively as responders, reviewers, and summarizers, striving to push the boundaries of its capabilities.
Appendix A Ethics Statement
This work adheres to the ICLR Code of Ethics, ensuring ethical compliance throughout all stages of the research. The MMRole-Data dataset was constructed using publicly available data, with rigorous quality control to prevent privacy risks. We acknowledge the potential biases present in the data and have taken proactive measures to ensure diversity and reduce these biases. Moreover, we recognize that MRPAs may generate dialogues that could misleadingly appear as actual statements made by real individuals. To prevent misunderstanding, it is essential to explicitly indicate that such content is simulated and does not represent genuine speech.
Appendix B Reproducibility Statement
For research reproducibility, the data, code, and models are all available at this GitHub repository. Additionally, the training settings of MMRole-Agent and the reward model in MMRole-Eval are presented in Section 6.2 and Section 6.4, respectively, while the training and inference code, along with the detailed assessment results of all evaluated MRPAs are submitted as supplementary materials. Moreover, the detailed prompts for dataset construction and performance evaluation are provided in Appendix E, Appendix I, Appendix G, and Appendix H, while the generated data is exemplified in Appendix C, Appendix D, and Figure 2.
Appendix C List of Characters
Figure 4: All character constructed in MMRole-Data, with the series to which the characters in the out-of-distribution test set belong being underlined.
Figure 4 lists all characters constructed in MMRole-Data, with the series to which the characters in the out-of-distribution test set belong being underlined.
Appendix D Examples of Character Profiles
@@
Appendix E Examples of the Two-Stage Generation Process for Hypothetical Real-Life Characters
@@
Figure 7 and Figure 8 present the prompts used to generate meta information and expand detailed profiles for hypothetical real-life characters, whereas Figure 9 exemplifies meta information for five randomly selected hypothetical real-life characters.
@@
Appendix F Manual Quality Control Strategies
For character profiles, we remove AI-assistant tones and unnecessary explanatory phrases, and reference reliable sources such as brainyquote.com to enhance the authenticity of catchphrases. Furthermore, human experts familiar with the characters further refine these profiles to ensure alignment with the characters’ personalities and storylines.
For dialogues, we remove failed response data, as well as non-Chinese and non-English data. Additionally, we eliminate content that replies in the tone of an AI assistant, meaningless modal words frequently output by GPT-4, action and scene descriptions, and unnecessary explanatory prefixes and suffixes.
Appendix G Prompts for Querying MRPAs
@@
Appendix H Prompts for GPT-4 Scoring and Reward Model Scoring
Appendix I Prompts for Dialogue Generation
@@
Appendix J RMSE and Pearson Results of the Reward Model
Table 9: The root mean squared error (RMSE) results. ‘Reward Model (GPT-4)’ denotes the scores evaluated by the reward model compared to those evaluated by GPT-4. ‘GPT-4 (humans)’ and ‘Reward Model (humans)’ signify the score gaps provided by GPT-4 and the reward model compared to the ground-truth score gaps provided by humans.
Evaluators (Ground Truth) IA Flu Coh ITR RA PC KC TC Overall
Reward Model (GPT-4) 0.1585 0.1076 0.1228 0.1334 0.1145 0.1564 0.1172 0.1778 0.1381
GPT-4 (humans) 0.1794 0.1421 0.1050 0.1253 0.1837 0.1826 0.1515 0.1946 0.1609
Reward Model (humans) 0.1356 0.1107 0.1465 0.1731 0.1810 0.2057 0.1793 0.2010 0.1695
Table 10: The Pearson correlation coefficient (Pearson) results. ‘Reward Model (GPT-4)’ denotes the scores evaluated by the reward model compared to those evaluated by GPT-4. ‘GPT-4 (humans)’ and ‘Reward Model (humans)’ signify the score gaps provided by GPT-4 and the reward model compared to the ground-truth score gaps provided by humans.
Evaluators (Ground Truth) IA Flu Coh ITR RA PC KC TC Overall
Reward Model (GPT-4) 0.7497 0.7344 0.7610 0.7955 0.8186 0.8167 0.8237 0.8129 0.8129
GPT-4 (humans) 0.6130 0.6736 0.9199 0.8184 0.7247 0.6997 0.7924 0.6985 0.7269
Reward Model (humans) 0.6561 0.3123 0.8033 0.8709 0.7321 0.7268 0.5832 0.5443 0.6502
As presented in Table 9 and Table 10, we report the root mean squared error (RMSE) and Pearson correlation coefficient (Pearson) results.
The overall RMSEs for Reward Model (GPT-4), GPT-4 (humans), and Reward Model (humans) are all relatively low, and those for GPT-4 (humans) and Reward Model (humans) are comparable, which are similar to the MAE results. Notably, the RMSE values are slightly higher than the MAE values, indicating some variability in the accuracy of both our reward model and GPT-4 across different test samples and evaluation metrics. This variability is expected, as the scoring difficulty varies across samples and metrics; for example, assessing personality consistency is significantly more complex than evaluating fluency.
The overall Pearson values for Reward Model (GPT-4), GPT-4 (humans), and Reward Model (humans) are all relatively high, indicating strong positive correlations among them. While the overall Pearson values for Reward Model (humans) are slightly lower than those for GPT-4 (humans), it performs well in metrics like Image-Text Relevance (0.8709), Response Accuracy (0.7321), and Personality Consistency (0.7268).
In summary, the combination of the MAE, RMSE, and Pearson results collectively demonstrates that our reward model effectively learns the evaluation abilities of GPT-4 and closely aligns with human evaluators.
Figure 14: The use cases of MMRole-Agent on the In-Test set.
Figure 15: The use cases of MMRole-Agent on the Out-Test set.
Figure 16: The use cases of GPT-4 on the In-Test set.
Figure 17: The use cases of GPT-4 on the Out-Test set.
Figure 18: The use cases of QWen-VL-Chat on the In-Test set.
Figure 19: The use cases of QWen-VL-Chat on the Out-Test set.
Appendix K Case Studies
Figures 14, 15, 16, 17, 18, and 19 present the use cases of our MMRole-Agent, GPT-4, and QWen-VL-Chat on both the In-Test and the Out-Test sets. Our observations indicate that both GPT-4 and MMRole-Agent perform strongly in multimodal role-playing, whereas Qwen-VL-Chat primarily functions as an AI assistant and struggles to adhere to role-playing instructions in inter-role dialogue scenarios. Moreover, we analyze the characteristics of MMRole-Agent from the following aspects:
  1. 1.
    Fundamental Conversational Skills: MMRole-Agent consistently fulfills the role-playing task by adhering closely to given instructions. Its outputs are not only fluent and coherent but also highly contextually appropriate.
  2. 2.
    Multimodal Understanding Abilities: MMRole-Agent produces outputs that maintain high relevance to visual inputs and effectively interpret image-based clues, even in complex multi-turn dialogues. Relevant examples are highlighted in purple and bold in the figures.
  3. 3.
    Role-Playing Qualities: MMRole-Agent effectively embodies the specified personality, tone, and experiences of its designated characters, showcasing distinctive speech patterns and ways of thinking. Relevant examples are highlighted in red and bold in the figures.
Appendix L Sensitivity Tests for MMRole-Agent on different Prompts
We conduct sensitivity tests on MMRole-Agent using different prompt templates. As shown in Table 11, we independently modify the system part and the character-designating part of the prompts. The performance of MMRole-Agent with these modified prompts remains nearly identical to that achieved with the original prompts (0.994). This indicates that MMRole-Agent is highly compatible with different prompt templates and does not exhibit signs of overfitting.
Table 11: The sensitivity test results for MMRole-Agent on different prompts.
Original Prompts Modified Prompts Overall
You are a dedicated role-playing assistant designed to immerse yourself fully in the character you are portraying. You are a highly skilled role-playing assistant, committed to fully immersing yourself in the character you embody. 0.995
Please step into the shoes of {role_name} from {role_series}. Imagine you are talking with a curious human about the given image. This requires a deep understanding of the character’s background, including their personality, experiences, abilities, and relationships. Imagine you are {role_name} from {role_series}, talking with a curious human about the given image. Draw on the character’s background, including their personality, experiences, abilities, and relationships. 0.996